OpenStudy (anonymous):

What is the surface area of the cone to the nearest whole number? Use 3.14 for π

OpenStudy (anonymous):

picture

OpenStudy (math&ing001):

Surface area = 3.14*16 (16+34)

OpenStudy (anonymous):

3.14 x 16 = 50.24 and 16+34= 50

OpenStudy (math&ing001):

This might be clearer: :) Surface area = 3.14*16*(16+34)

OpenStudy (math&ing001):

Here's how I got it :

OpenStudy (anonymous):

i still dont get it :/

OpenStudy (anonymous):

it gave me 7626.62

OpenStudy (anonymous):

but its not in my answer choices

OpenStudy (math&ing001):

Ok give me your answer choices :)

OpenStudy (math&ing001):

And t's not 7626.62

OpenStudy (anonymous):

this are my answer choices : A.) 1,708 m2 B.) 3,316 m2 C.) 2,512 m2 D.) 4,220 m2

OpenStudy (math&ing001):

You told me before, you got 3.14 x 16 = 50.24 and 16+34= 50. Can you multiply them ? 50*50.24

OpenStudy (anonymous):

2512 so the answer is C

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

thanks :)

OpenStudy (math&ing001):

Welcome ^^